OpenStudy (anonymous):

What is 3/12 as a decimal? Please answer I will reward! ♥ ☻ ♥

OpenStudy (anonymous):

3/12 = 1/4 1/4 = 25/100 0.25

OpenStudy (anonymous):

Thx!